OpenStudy (anonymous):

The product of (a − b)(a − b) is a perfect square trinomial. Sometime, always, or never?

OpenStudy (anonymous):

yes, always.

OpenStudy (anonymous):

ok thank you! but why?

OpenStudy (anonymous):

note that (a-b)(a-b) = (a-b)^2

OpenStudy (anonymous):

and (a-b)^2 leads to a perfect square trinomial. :))

OpenStudy (anonymous):

by foiling (a-b)^2 = a^2 -2ab +b^2 , which is a perfect square trinomial.

OpenStudy (anonymous):

ohhh i see ok well thank you :)

OpenStudy (anonymous):

most welcome, good day!